OpenStudy (anonymous):

The number 912 is evenly divisible by which of the following sets of numbers? A. 2, 6, 9 B. 2, 3, 6 C. 2, 3, 6, 9 D. 2, 3, 9

OpenStudy (lgbasallote):

have you tried trial and error?

OpenStudy (lasttccasey):

912 is not divisible by 9 so B is the only one left. Just "plug n chug" to see what works.\[912/2=456\]\[912/3=304\]\[912/6=152\]\[912/9=101.33333333333333333333333333333 \]Which ones don't have decimals?

OpenStudy (anonymous):

912/9

OpenStudy (lasttccasey):

Right, 9 does not evenly divide into 912 so the others are your correct answers, 2, 3, and 6.

OpenStudy (anonymous):

well the no. 912 consists of 19.16.3 So we can Apply 1(3) only !